What is the speed of light in benzene?
1
Expert's answer
2020-11-16T07:52:17-0500

The refractive index of benzene is 1.4957.

The speed of light in benzene is


"c_b=\frac cn=\frac{1}{n\sqrt{\mu_0\epsilon_0}}=200436222.5\text{ m/s}."